if a number is added to 3 more than 5 times itself the result is 30

Found 2 solutions by algebrahouse.com, vleith:
Answer by algebrahouse.com(1079)

"If a number is added to 3 more than 5 times itself the result is 30"

x = the number

x + 5x + 3 = 30 {a number added to 5 more than three times itself equals 30}
6x + 3 = 30 {combined like terms}
6x = 27 {subtracted 3 from both sides}
x = 27/6 {divided both sides by 6}
= 9/2 {reduced}
= 4 1/2 {wrote as a mixed number}
www.algebrahouse.com

Answer by vleith(2825)